### v1

# Temporal Smoothness-Aware Rate-Distortion Optimized 4D Gaussian Splatting

Abstract
Dynamic 4D Gaussian Splatting (4DGS) effectively extends the high-speed rendering capabilities of 3D Gaussian Splatting (3DGS) to represent volumetric videos. However, the large number of Gaussians, substantial temporal redundancies, and especially the absence of an entropy-aware compression framework result in large storage requirements. Consequently, this poses significant challenges for practical deployment, efficient edge-device processing, and data transmission. In this paper, we introduce a novel end-to-end RD-optimized compression framework tailored for 4DGS, aiming to enable flexible, high-fidelity rendering across varied computational platforms. Leveraging Fully Explicit Dynamic Gaussian Splatting (Ex4DGS), one of the state-of-the-art 4DGS methods, as our baseline, we start from the existing 3DGS compression methods for compatibility while effectively addressing additional challenges introduced by the temporal axis. In particular, instead of storing motion trajectories independently per point, we employ a wavelet transform to reflect the real-world smoothness prior, significantly enhancing storage efficiency. This approach yields significantly improved compression ratios and provides a user-controlled balance between compression efficiency and rendering quality. Extensive experiments demonstrate the effectiveness of our method, achieving up to 91$\times$ compression compared to the original Ex4DGS model while maintaining high visual fidelity. These results highlight the applicability of our framework for real-time dynamic scene rendering in diverse scenarios, from resource-constrained edge devices to high-performance environments.
1 Introduction
The advent of 3D Gaussian Splatting (3DGS) 3dgs marked a significant step in real-time radiance field rendering, offering superior speed and simpler training compared to Neural Radiance Field (NeRF)-based approaches nerf ; nerfies ; wu2023efficient ; dihlmann2024signerf ; zhong2024structured ; ren2024nerf . Furthermore, the Gaussian Splatting framework is not restricted to static scenes but has been effectively extended to represent dynamic, or 4D, scenes such as videos by capturing temporal variations 4dgs ; yang2023real ; yang2024deformable ; duisterhof2023deformgs ; duisterhof2023md ; girish2024queen .
Dynamic 4D Gaussian Splatting (4DGS) methods typically require a substantial number of Gaussians and extensive Spherical Harmonics (SH) parameters to represent volumetric video accurately. Additionally, dynamic points in these methods are often stored independently per timestamp, inherently causing significant temporal redundancies. Consequently, such characteristics lead to large models and excessive storage demands, severely restricting applicability in resource-constrained scenarios, including edge computing and efficient data transmission. Although several existing techniques liu2025light4gs ; yuan20251000+ ; zhang2024mega ; liu2024lgs have explored model size reduction, there is currently no entropy-aware, bit-level compression method for 4DGS that enables flexible rate-distortion optimization.
In this paper, we propose a novel rate-distortion (RD) optimized compression method specifically designed for 4D Gaussian Splatting (4DGS), enabling real-time applications across diverse computing environments, ranging from edge devices to general computing platforms. We adopt Fully Explicit Dynamic Gaussian Splatting (Ex4DGS) ex4dgs as our baseline model and introduce a novel approach that leverages inherent video characteristics beyond traditional RD-optimization. Specifically, following Ex4DGS, we decompose the scene into static and dynamic components. For components common to both 3DGS and 4DGS, we apply masking and vector quantization after model training, following the approach of Wang et al. wang2024rdo3dgs . Although Ex4DGS reduces redundancy by storing dynamic points only at selected keyframes instead of every frame, the positions at these keyframes are stored independently, inherently limiting the exploitation of temporal redundancy and thus reducing compression efficiency. Instead of utilizing the compression method for 3DGS directly, we leverage the prior that each point moves along a smooth trajectory. To this end, we apply a novel wavelet transform to each point’s trajectory and discard high-frequency detail components, significantly reducing storage without severely compromising fidelity. Additionally, we conduct a fine-grained analysis, evaluating the rate-distortion impact of quantizing individual dynamic opacity elements. Based on this analysis, we selectively apply quantization only to those elements that offer a favorable trade-off between compression rate and rendering quality.
Our method achieves model sizes as low as only 1.1% on average of the original Ex4DGS model while maintaining reasonable visual fidelity. Furthermore, our framework allows users to flexibly control the trade-off between distortion and data rate based on specific scenario requirements.
Our contributions can be summarized as follows:
  • We propose a novel end-to-end rate-distortion optimized compression framework specifically designed for 4D Gaussian Splatting representations, using Ex4DGS as the baseline.
  • To the best of our knowledge, this work is the first to introduce a bit-level rate-distortion optimizable compression framework for dynamic 4D Gaussian representations.
  • We apply the pointwise wavelet transform to further compress the position trajectories of dynamic Gaussian points, effectively leveraging temporal redundancy inherent in 4D representations.
  • Our method achieves significant compression ratios (up to 91$\times$) for dynamic scenes while maintaining reasonable visual fidelity and offering flexible control over the rate-distortion trade-off suitable for diverse computing platforms.
2 Related Work
2.1 Neural Rendering
Neural Radiance Fields (NeRF) introduced a novel view synthesis approach by learning density and color via a multi-layer perceptron from 3D coordinates and viewing directions nerf . Despite their effectiveness, NeRF-based methods suffer from slow rendering and training speed. Follow-up works accelerated rendering by employing multi-resolution hash grids instantngp or optimizing sparse voxel grids directly fridovich2022plenoxels , though often at the expense of increased memory usage.
3D Gaussian Splatting (3DGS) 3dgs emerged as a fast and memory-efficient alternative by representing scenes with anisotropic 3D Gaussians rather than dense voxels or networks. By optimizing the parameters, including mean, covariance, and color, 3DGS achieves real-time rendering at high resolutions using efficient algorithms instead of volumetric ray marching. Recent studies demonstrate that 3DGS matches or surpasses NeRF in rendering quality, speed, and scalability.
Dynamic scene modeling extends this idea into 4D Gaussian Splatting (4DGS), incorporating time as an additional dimension. Dynamic NeRF methods nerfies ; hypernerf learn deformation fields or multi-space embeddings to handle scene motion, but they remain computationally heavy and slow to render. In contrast, 4DGS 4dgs ; yang2024deformable was introduced as a unified representation for spatiotemporal data, using a set of 4D Gaussians with per-Gaussian motion trajectories or a deformation network to render moving scenes in real time. Overall, Gaussian Splatting has rapidly become a promising alternative to NeRF, offering comparable quality with significantly improved rendering speed, which is crucial for interactive applications.
2.2 Gaussian Splatting Compression
Explicit scene representations like 3D Gaussian Splatting (3DGS) 3dgs enable real-time, high-fidelity rendering but inherently introduce substantial memory requirements. These arise from the need to store numerous Gaussian primitives, each characterized by parameters for position, shape, opacity, and color (often represented by complex spherical harmonics). This significant memory footprint has spurred the development of various compression techniques specifically designed for 3DGS.
Recent approaches tailored for compressing 3DGS include Compact3DGS compact3dgs , which employs trainable masking to prune redundant Gaussians, CompGS navaneet2024compgs , adopting vector quantization to store Gaussians, HAC++hacpp , leveraging hash-grid-assisted entropy coding, and Context 3DGS wang2024contextgs , utilizing an autoregressive, anchor-based context model for hierarchical Gaussian compression.
Extending these techniques to dynamic scenes poses greater memory challenges due to temporal variations. Recent specialized solutions include Light4GS liu2025light4gs , combining spatio-temporal pruning with entropy coding; 4DGS-1K yuan20251000+ , optimizing Gaussian reuse and visibility masks for ultra-fast rendering; MEGA zhang2024mega , replacing spherical harmonics with lightweight predictors for significant compression; and LGS liu2024lgs , pruning motion-insensitive Gaussians and redundant deformation fields. These methods highlight the growing focus on efficient and high-quality rendering of dynamic scenes. QUEEN girish2024queen introduces integer-based quantization enabling entropy coding; however, it does not explicitly optimize entropy within a rate-distortion framework, leaving rate-distortion optimization for dynamic Gaussian Splatting unexplored.
2.3 End-to-End Rate-Distortion Optimized Compression
Rate-distortion (RD) optimization balances bitrate against reconstruction quality, a core principle extensively used in traditional and neural compression methods. End-to-end image compression frameworks by Ballé et al. balle2017end introduced autoencoder-based models jointly trained to minimize distortion and bitrate, laying the foundation for differentiable compression. Follow-up studies improved efficiency through advanced context modeling and latent priors cheng2020learned ; li2021deep ; he2022elic .
Recently, RD-optimization concepts have been applied to neural scene representations. For example, Takikawa et al. presented a neural field that continuously adjusts detail level and bitrate via a tunable parameter takikawa2022variable . Specifically for 3D Gaussian Splatting, Wang et al. wang2024rdo3dgs proposed an RD-optimization-based framework that jointly optimizes Gaussian parameters with learnable pruning and entropy-constrained quantization, achieving substantial reductions in model size with minimal quality loss. Such approaches highlight the benefit of integrating RD-optimization directly into the training process, enabling flexible compression tailored for diverse computational environments. While compression techniques exist for 4DGS and RD-optimization has been applied to 3DGS, a dedicated end-to-end RD-optimized framework specifically for 4DGS, offering flexible control over the rate-quality trade-off by effectively exploiting spatiotemporal structure, is still lacking. In this work, we develop a novel method to apply end-to-end RD-optimized compression to the 4D domain, which is related to spatiotemporal representation.
3 Method
3.1 Preliminary
In this section, we first introduce the fundamental concepts behind 3D Gaussian Splatting (3DGS) 3dgs . We then briefly describe its extension to 4D scenes and detail our baseline model, Ex4DGS ex4dgs , which explicitly models temporal dynamics. Finally, we outline the principles of end-to-end rate-distortion optimization as applied to Gaussian Splatting.
3D Gaussian Splatting. 3D Gaussian Splatting (3DGS) 3dgs represents scenes using explicit 3D Gaussian primitives. Each primitive $i$ is parameterized by a position (mean) $\boldsymbol{\mu}_{i}\in\mathbb{R}^{3}$, a covariance matrix $\boldsymbol{\Sigma}_{i}\in\mathbb{R}^{3\times 3}$, an opacity value $\alpha_{i}\in[0,1]$, and view-dependent color represented by Spherical Harmonics (SH) coefficients $\mathbf{c}_{i}$. The 3D Gaussian function is defined as:
$G(\mathbf{x})=e^{-\frac{1}{2}(\mathbf{x}-\boldsymbol{\mu}_{i})^{T}\boldsymbol{\Sigma}_{i}^{-1}(\mathbf{x}-\boldsymbol{\mu}_{i})}\ .$ (1)
The 3D covariance matrix $\boldsymbol{\Sigma}_{i}$ defines the shape and orientation and is efficiently parameterized using a scaling vector $\mathbf{s}_{i}\in\mathbb{R}^{3}$ and a rotation quaternion $\mathbf{q}_{i}\in\mathbb{R}^{4}$, such that $\boldsymbol{\Sigma}_{i}=\mathbf{R}_{i}\mathbf{S}_{i}\mathbf{S}_{i}^{T}\mathbf{R}_{i}^{T}$, where $\mathbf{R}_{i}$ is the rotation matrix derived from $\mathbf{q}_{i}$ and $\mathbf{S}_{i}$ is a diagonal scaling matrix derived from $\mathbf{s}_{i}$. To render an image, these 3D Gaussians are projected onto the 2D image plane and then “splatted”. The final pixel color $\mathbf{C}$ is computed by alpha-blending the contributions from $N$ Gaussians sorted by depth:
$\mathbf{C}=\sum_{i=1}^{N}\mathbf{c}_{i}\alpha_{i}\prod_{j=1}^{i-1}(1-\alpha_{j})\ ,$ (2)
where $\alpha_{i}$ here denotes the projected 2D Gaussian’s opacity, derived from $\alpha_{i}$ and the 2D covariance. This differentiable rendering process allows for end-to-end optimization of all Gaussian parameters ($\boldsymbol{\mu}_{i},\mathbf{s}_{i},\mathbf{q}_{i},\alpha_{i},\mathbf{c}_{i}$) using gradient descent, typically initialized from a sparse point cloud obtained via Structure-from-Motion (SfM). 3DGS employs adaptive density control, involving splitting, cloning, and pruning Gaussians during optimization.
4D Gaussian Splatting. 4D Gaussian Splatting (4DGS) 4dgs adapts 3DGS for dynamic scenes. We build upon Ex4DGS ex4dgs , which separates Gaussians into static ($\mathcal{G}_{s}$) and dynamic ($\mathcal{G}_{d}$) components. Static Gaussians ($\mathcal{G}_{s}$) maintain constant covariance ($\boldsymbol{\Sigma}_{s}$), opacity ($\alpha_{s}$), and SH coefficients ($\mathbf{c}_{s}$), with only position $\boldsymbol{\mu}_{s}(t)$ changing linearly over the scene duration $L$:
$\boldsymbol{\mu}_{s}(t)=\boldsymbol{\mu}_{0}+\frac{t}{L}\boldsymbol{\mu}_{disp}$ (3)
where $\boldsymbol{\mu}_{0}$ is a pivot position and $\boldsymbol{\mu}_{disp}$ is the learned translation. Dynamic Gaussians ($\mathcal{G}_{d}$) use a keyframe approach, storing parameters only at sparse, uniformly sampled keyframes $\mathcal{K}=\{t\mid t=nI,n\in\mathbb{Z},t\in\mathcal{T}\}$. Properties at intermediate times $t$ are interpolated: position $\boldsymbol{\mu}_{d}(t)$ using Cubic Hermite Interpolation (CHip), quaternion rotation $\mathbf{q}_{d}(t)$ using Spherical Linear Interpolation (Slerp), and temporal opacity $\sigma_{t}(t)$ using a custom two-Gaussian mixture model (parameterized by means $a_{s}^{o},a_{f}^{o}$ and variances $b_{s}^{o},b_{f}^{o}$) to handle appearance/disappearance. For dynamic Gaussians in Ex4DGS, scale ($\mathbf{s}_{d}$) and SH coefficients ($\mathbf{c}_{d}$) are typically assumed constant over time.
3.2 4D Dynamic Gaussian Compression
Building upon the Ex4DGS framework ex4dgs , we introduce an end-to-end compression framework specifically tailored for dynamic 4D Gaussian representations. We follow the general principles established by Wang et al. wang2024rdo3dgs for compatibility with 3DGS compression, extending them to accommodate the unique challenges presented by dynamic scenes.
Gaussian and SH Coefficient Masking. To reduce model complexity, redundant Gaussians are pruned using learnable masks $\phi_{i}$. A soft mask $\phi_{i}^{soft}=\text{sigmoid}(\phi_{i})$ is generated, then binarized to a hard mask $\phi_{i}^{hard}\in\{0,1\}$ using a threshold $\phi_{th}$ and the straight-through estimator (STE): $\phi_{i}^{hard}=\text{sg}(\mathbb{I}(\phi_{i}^{soft}>\phi_{th})-\phi_{i}^{soft})+\phi_{i}^{soft}$. This hard mask is applied to key attributes like scale $s_{i}$ and opacity $\alpha_{i}$. The associated pruning cost, contributing to the rate term $\mathcal{L}_{rate}$, is:
$\mathcal{L}_{GSprune}=\frac{1}{N}\sum_{i}\phi_{i}^{soft}$ (4)
Similarly, redundant SH coefficients are adaptively pruned using learnable masks $\theta_{i}^{(l)}$ per Gaussian $i$ and degree $l\geq 1$. Soft masks $\theta_{i}^{(l)soft}=\text{sigmoid}(\theta_{i}^{(l)})$ are binarized to hard masks $\theta_{i}^{(l)hard}$ via STE using a threshold $\theta_{th}$, too. The hard mask is applied to zero out corresponding coefficients $\hat{c}_{i}^{(l)}=\theta^{(l)hard}c_{i}^{(l)}$. The weighted pruning loss contributing to $\mathcal{L}_{rate}$ is:
$\mathcal{L}_{SHprune}=\frac{1}{N}\sum_{i}\sum_{l=1}^{k}\frac{2l+1}{(k+1)^{2}-1}\theta_{i}^{(l)soft}$ (5)
This masking strategy is uniformly applied to both static and dynamic Gaussians, and experiments confirmed effective pruning across both categories. The results are available in Appendix C.
SH Coefficient, Rotation, and Scale Compression. Spherical Harmonics (SH) coefficients, rotation, and scale parameters exist for both static and dynamic Gaussians. These parameters are compressed using entropy-constrained vector quantization (ECVQ) ecvq , following Wang et al. wang2024rdo3dgs . Recognizing potential differences in compression sensitivity, scalar quantization is selectively applied.
Specifically, rotation parameters for dynamic Gaussians differ from static Gaussians in that they are defined at each timestamp. Although storing residual rotations was initially considered, preliminary experiments indicated that the entropy gain was marginal compared to the complexity of maintaining additional codebooks. Hence, we opted to store all rotation parameters directly.
Opacity. Unlike traditional 3D Gaussian methods, Ex4DGS parameterizes dynamic opacity temporally by defining a scale function with center parameters ($a_{s}^{o},a_{f}^{o}$) and variance parameters ($b_{s}^{o},b_{f}^{o}$), effectively acting as an on/off temporal mask multiplied with a base opacity ($\alpha^{d}$). Given that the base opacity for dynamic points shares similar properties with the static opacity ($\alpha^{s}$), we follow Wang et al. wang2024rdo3dgs and apply scalar quantization to the base opacity. However, the temporal center and variance parameters, being expressed in time units rather than opacity values, exhibit different sensitivities to quantization. Empirical tests indicated that variance parameters ($b_{s}^{o},b_{f}^{o}$) were particularly sensitive to quantization artifacts, hence, we avoid aggressive quantization for these parameters to maintain visual fidelity. Detailed empirical validation of this decision is presented in Section 4.3.
3.3 Wavelet Compression for Dynamic Positions
To efficiently represent the motion trajectories of dynamic Gaussians ($\boldsymbol{\mu}_{d}(t)\in\mathbb{R}^{T\times 3}$), we employ a single-level Haar wavelet transform along the temporal dimension. Formally, given the dynamic positions $\boldsymbol{\mu}_{d}=[\mathbf{p}_{1},\mathbf{p}_{2},\dots,\mathbf{p}_{T}]^{\top}$ with $\mathbf{p}_{t}\in\mathbb{R}^{3}$, the wavelet transform is applied as:
$\mathbf{F}=\mathbf{W}\boldsymbol{\mu}_{d},\quad\mathbf{F}=\begin{bmatrix}\mathbf{F}_{a}\ \mathbf{F}_{d}\end{bmatrix}$ (6)
Here, $\mathbf{W}$ denotes the Haar wavelet transform matrix, decomposing the trajectory into approximation (low-frequency) coefficients $\mathbf{F}_{a}\in\mathbb{R}^{\frac{T}{2}\times 3}$ and detail (high-frequency) coefficients $\mathbf{F}_{d}\in\mathbb{R}^{\frac{T}{2}\times 3}$. Recognizing that coarse motion patterns captured by $\mathbf{F}_{a}$ are perceptually more significant, we explicitly discard the detail coefficients $\mathbf{F}_{d}$ by zeroing them out. And the compressed trajectory $\hat{\boldsymbol{\mu}}_{d}$ is then reconstructed via the inverse Haar transform, which is equivalent to multiplying by the transpose of $\mathbf{W}$ due to its orthogonality:
$\mathbf{F}_{\text{masked}}=\begin{bmatrix}\mathbf{F}_{a}\ \mathbf{0}\end{bmatrix},\ \ \hat{\boldsymbol{\mu}}_{d}=\mathbf{W}^{\top}\mathbf{F}_{\text{masked}}$ (7)
This approach significantly reduces storage and transmission requirements while effectively preserving essential dynamic motion information.
3.4 End-to-End Rate-Distortion Optimization
Following Bellé et al. balle2017end , our framework performs end-to-end rate-distortion optimization. Our framework jointly optimizes all relevant parameters by minimizing a composite loss function. These parameters include the original Ex4DGS parameters, those from our differentiated quantization (such as codebooks and entropy models for ECVQ), and the wavelet-compressed dynamic position representations. This total loss, $\mathcal{L}_{total}$, integrates the standard rate-distortion objective with regularization terms inherent to Ex4DGS:
$\mathcal{L}_{total}=\mathcal{L}_{dist}+\lambda_{R}\mathcal{L}_{rate}+\lambda_{reg}\mathcal{L}_{reg}$ (8)
Here, $\mathcal{L}_{dist}$ represents the reconstruction fidelity loss, typically combining L1 loss and D-SSIM between the rendered image $\mathbf{C}$ and the ground truth image $\mathbf{C}_{GT}$:
$\mathcal{L}_{dist}=(1.0-\lambda_{dssim})\mathcal{L}_{L1}(\mathbf{C},\mathbf{C}_{GT})+\lambda_{dssim}(1.0-\text{SSIM}(\mathbf{C},\mathbf{C}_{GT}))$ (9)
The term $\mathcal{L}_{rate}$ quantifies the estimated bitrate cost of the compressed representation. It aggregates several components aimed at reducing model size. These include entropy-based costs derived from quantized attribute representations (e.g. VQ indices) and potentially auxiliary losses that promote sparsity within the model, such as terms encouraging the pruning of redundant Spherical Harmonics coefficients or entire Gaussian primitives. These components are collectively balanced by the rate weighting factor $\lambda_{R}$. When ECVQ is employed, the objective intrinsically balances the rate cost, and the quantization distortion for the respective attributes during codeword selection and codebook learning. Furthermore, $\mathcal{L}_{reg}$ incorporates regularization terms adapted from Ex4DGS ex4dgs to encourage plausible scene dynamics. This includes terms penalizing excessive linear displacement for static points and terms ensuring temporal smoothness in the motion and rotation trajectories defined by the dynamic keyframes.
By optimizing this comprehensive loss function end-to-end, our framework learns the compression mechanisms (quantization, pruning, wavelet processing via its impact on $\mathcal{L}_{dist}$ and $\mathcal{L}_{rate}$) concurrently with the scene representation itself. This joint process directly balances the bitrate costs against the final rendering quality and dynamic consistency, guided by the various lambda hyperparameters.
4 Experiment
4.1 Experimental Settings
Datasets. To validate the effectiveness of the proposed method, we conduct various experiments on two real-world datasets, including Neural 3D Video (N3V) li2022n3v and Technicolor sabater2017techni . Because the proposed method is built upon Ex4DGS ex4dgs , we match the evaluation protocol, it is originally adopted by the previous work attal2023hyperreel . Specifically, we use the entire set of N3V and the five scenes of Technicolor (Birthday, Fabien, Painter, Theater, and Train).
Metrics. As in the previous works, we report standard metrics including Peak Signal-to-Noise Ratio (PSNR), Structural Similarity Index Measure (SSIM), and Learned Perceptual Image Patch Similarity (LPIPS) lpips for quantitative evaluation. We also measure the storage size for each compression level to evaluate if the proposed method achieves reasonable performance across different compression strengths. By varying the compression parameters, we generate rate-distortion (RD) curves to comprehensively analyze the trade-off between storage size (rate) and visual quality (distortion). Additionally, we report frames-per-second (FPS) to measure the rendering speed.
Implementation Details. We first train the base model using Ex4DGS ex4dgs until densification terminates. Subsequently, we apply our proposed RD-optimized method, incorporating Gaussian point pruning, adaptive spherical harmonics (SH) pruning, entropy-constrained vector quantization (ECVQ), and the novel pointwise wavelet transform. To systematically analyze performance at varying compression strengths, we define six compression levels (Levels 1–6), adjusting pruning hyperparameters as follows: for Gaussian point pruning, $\lambda_{GSprune}$ is set to [0.05, 0.02, 0.01, 0.005, 0.002, 0.0005]; for SH coefficient pruning, $\lambda_{SHprune}$ is set to [0.5, 0.2, 0.1, 0.05, 0.02, 0.005], respectively. All experiments are conducted on an NVIDIA RTX 3090 GPU.
Figure 1: Average RD performance comparison between our method (Levels 1-6) and the Ex4DGS ex4dgs on the Neural 3D Video (N3V) dataset. Per-scene details are provided in Appendix A.
Figure 2: Average RD performance comparison between our method (Levels 1-6) and the Ex4DGS ex4dgs on the Technicolor dataset. Per-scene details are provided in Appendix A.
Figure 3: Rate-distortion impact of using wavelet transform for dynamic positions on the N3V dataset. We compare the proposed method (‘Ours’, with wavelet transform) against the same model without wavelet transform (‘w/o Wavelet Transform’). Axes show avg. PSNR (dB) vs. avg. model size (MB).
Figure 4: Qualitative results on the Technicolor dataset. We compare Ground Truth, Ex4DGS ex4dgs , and ours at compression levels 6 and 1. PSNR (dB) / Size (MB) are shown below each image.
Figure 5: Qualitative results on the N3V dataset. We compare Ground Truth, Ex4DGS ex4dgs , and ours at compression levels 6 and 1. PSNR (dB) / Size (MB) are shown below each image.
4.2 Rate-Distortion Performance
In this section, we analyze the end-to-end rate-distortion (RD) optimized compression performance of the proposed method. Our primary goal is to leverage the spatio-temporal characteristics of videos to achieve compression efficiency beyond the limits of existing methods, while offering flexible control over the rate-distortion trade-off. This goes beyond simple masking and quantization by incorporating video-specific properties. To achieve different operating points on the RD curve, we generate compressed Gaussians of varying sizes by adjusting the strength of Spherical Harmonics (SH) and Gaussian point pruning and quantization, similar to RD3DGS wang2024rdo3dgs . This is primarily controlled by adjusting the weight $\lambda$ for $\mathcal{L}_{GSprune}$ and $\mathcal{L}_{SHprune}$. Figures 1 and 2 present the rate-distortion curves (PSNR, SSIM, LPIPS vs. Model size in MB) for the N3V dataset li2022n3v and the Technicolor dataset sabater2017techni , respectively. Due to the space constraint, we provide the per-scene results in Table 4. Furthermore, Figures 5 and 4 provide qualitative
Table 1: Ablation on quantizing opacity components ($\alpha^{s}$, $\alpha^{d}$, $a_{*}^{o}$, $b_{*}^{o}$) for compression levels 1 and 6 on the N3V dataset. PSNR (dB) and relative model size reduction (%) are reported. $\checkmark$ indicates the component is quantized.
$\alpha^{s}$ $\alpha^{d}$ $a_{*}^{o}$ $b_{*}^{o}$ $\mathbb{L}1$ $\mathbb{L}6$
- - - - 26.86 (0.00%) 29.57 (0.00%)
$\checkmark$ 26.86 (4.16%) 29.57 (4.01%)
$\checkmark$ $\checkmark$ 26.86 (3.84%) 29.57 (3.39%)
$\checkmark$ $\checkmark$ $\checkmark$ 26.86 (4.97%) 29.57 (5.30%)
$\checkmark$ $\checkmark$ $\checkmark$ 26.17 (5.16%) 28.52 (5.58%)
$\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 26.17 (5.98%) 28.52 (6.87%)
comparisons, showcasing the visual fidelity of our method at key compression levels against original Ex4DGS ex4dgs . The results demonstrate that our proposed approach, incorporating techniques like differentiated quantization and the wavelet transform, achieves significant model size reduction (e.g., up to 91$\times$ compression) compared to Ex4DGS while maintaining reasonable visual quality. This validates the effectiveness of our method in providing a flexible trade-off between compression rate and rendering quality suitable for diverse computational requirements.
4.3 Ablation Study
Opacity Quantization. We validate the opacity quantization strategy described in Section 3.2 through ablation experiments summarized in Table 1. Specifically, we examine the quantization sensitivity of four parameters: static opacity ($\alpha^{s}$), dynamic opacity ($\alpha^{d}$), center parameters ($a_{s}^{o},a_{f}^{o}$), and variance parameters ($b_{s}^{o},b_{f}^{o}$). Although static and dynamic opacity share similar characteristics, we evaluate them separately, considering the potential heightened sensitivity of dynamic opacity. The results demonstrate that quantizing additional opacity-related parameters consistently improves storage efficiency. However, notably in the last two rows, quantizing variance parameters significantly reduces PSNR despite achieving marginal additional storage benefits. Based on the observation, we conclude that quantization should be limited to the static and dynamic opacity and center parameters to maintain a balanced trade-off between rendering fidelity and compression efficiency.
Wavelet Transform for Dynamic Positions. We further evaluate the effectiveness of wavelet-based compression applied specifically to the dynamic position trajectories ($\boldsymbol{\mu}_{d}(t)$). As detailed in Section 3, our proposed method leverages a Haar wavelet transform to capture low-frequency motion components while explicitly discarding high-frequency details. Figure 3 compares the rate-distortion performance of our method with wavelet transform against the baseline without it, clearly highlighting the benefit of incorporating wavelet compression. Notably, employing the wavelet transform consistently improves performance across all tested compression levels, not only reducing the storage size but also enhancing rendering fidelity. Specifically, at compression level 1, our method achieves approximately 19% storage reduction alongside a PSNR gain of 0.19 dB. Even at the highest fidelity setting (level 6), we observe around 26% storage savings coupled with a PSNR improvement of 0.09 dB. These results emphasize that exploiting the inherent smoothness of point trajectories through wavelet-based compression not only effectively reduces redundancy in dynamic positions but also enhances trajectory modeling accuracy, leading to improved overall RD performance.
4.4 Comparative Evaluation
Table 2: Quantitative results on the N3V dataset comparing PSNR (dB), average model size (MB), and FPS. Our proposed RD-optimized method offers flexible compression levels. Level 6 achieves a significant size reduction (e.g., about 10$\times$ reduction to 11.06MB) compared to the Ex4DGS ex4dgs while maintaining competitive quality. Level 1 demonstrates the capability for aggressive compression (about 90$\times$ size reduction to 1.26MB) for highly resource-constrained scenarios.
[…]
Coffee
Martini
Cook
Spinach
Cut Roasted
Beef
Flame
Salmon
Flame
 Steak
Sear
 Steak
Average
Size
FPS
NeRF-based
NeRFPlayer song2023nerfplayer 31.53 30.56 29.35 31.65 31.93 29.13 30.69 5130 0.05
HyperReel attal2023hyperreel 28.37 32.30 32.92 28.26 32.20 32.57 31.10 360 2
K-Plane fridovich2023k 29.99 32.60 31.82 30.44 32.38 32.52 31.63 311 0.3
MixVoxels-L wang2023mixed 29.63 32.25 32.40 29.81 31.83 32.10 31.34 500 37.7
MixVoxels-X wang2023mixed 30.39 32.31 32.63 30.60 32.10 32.33 31.73 500 4.6
Gaussian Splatting-based
4DGS 4dgs 26.51 32.11 31.74 26.93 31.44 32.42 30.19 6057 72.0
4DGaussians yang2023real 26.69 31.89 25.88 27.54 28.07 31.73 28.63 34 146.6
3DGStream sun20243dgstream 27.75 33.31 33.21 28.42 34.30 33.01 31.67 1200 -
Ex4DGS ex4dgs 28.79 33.23 33.73 29.29 33.91 33.69 32.11 115 72.3
Ours [$\mathbb{L}$1] 24.97 28.38 28.02 24.16 28.33 28.39 27.04 1.26 (-98.9%) 163.0
Ours [$\mathbb{L}$6] 26.17 30.69 31.45 26.58 31.34 31.70 29.66 11.06 (-90.4%) 100.9
This section evaluates the proposed method against existing models. Our primary goal is a framework for flexible rate-distortion trade-offs in dynamic 4D scene representation, not just state-of-the-art rendering. The comparison highlights the method’s quality-size characteristics and its ability to significantly compress data while maintaining reasonable fidelity.
We report the PSNR, storage size, and FPS for the N3V dataset in Table 2. The most notable advantage of our method is its significant size reduction. At compression Level 1, our model achieves a remarkable 98.9% size reduction compared to the baseline Ex4DGS, resulting in an average size of only 1.26 MB. Even when compared to 4DGaussians, the previously most storage-efficient method, our approach demonstrates over 25 times greater efficiency at the same compression level. Our method significantly outperforms existing models in rendering speed, reaching up to 163 FPS at Level 1, suitable for real-time applications.
Table 3: Quantitative results averaged over the Technicolor dataset. We compare the proposed method (L1, L6) against the Ex4DGS ex4dgs using standard quality metrics and model size.
PSNR SSIM LPIPS Size(MB) FPS(s)
Ex4DGS ex4dgs 33.62 0.916 0.088 140.2 72.3
Ours [$\mathbb{L}$6] 32.20 0.904 0.100 19.6(-86.0%) 113.1
Ours [$\mathbb{L}$1] 28.60 0.822 0.232 2.1(-98.5%) 213.9
While there is an expected trade-off in rendering quality (PSNR) due to aggressive compression, our model still produces results that are sufficiently competitive. Particularly noteworthy is the comparison at Level 6, where our method not only has a smaller model size (11.06 MB) but also a higher average PSNR (29.66 dB) compared to 4DGaussians (34 MB, 28.63 dB).
5 Discussion
Conclusion. We presented the first end-to-end rate-distortion (RD) optimized compression framework for 4D Gaussian Splatting (4DGS) to address the challenge of large model sizes hindering deployment. Leveraging Ex4DGS, our method uses static/dynamic decomposition with adaptive RD-guided quantization and pointwise wavelet transform to achieve flexible compression. Our approach demonstrates significant model size reduction (down to approximately 1.1% of the original size ) while maintaining reasonable visual quality across standard benchmarks and offering user control over the rate-quality trade-off. This significantly improves the practicality of 4DGS for real-time rendering on diverse computational platforms, including resource-constrained devices.
Limitation. We observed motion blur artifacts of dynamic objects at high compression levels. Additionally, dynamic points still account for a substantial portion of total storage, highlighting significant room for improvement in compressing dynamic components (Appendices D and E). Future work includes refining the loss function, for example by applying separate masking weights for static and dynamic points, to further enhance dynamic point compression and overall visual quality.
Broader Impacts. Our work significantly simplifies the storage, transmission, and deployment of volumetric videos, directly contributing to the widespread adoption and accessibility of volumetric content. In the long term, we anticipate that this research can serve as a foundational step toward establishing standards for volumetric video formats. Nevertheless, easier distribution may also enable rapid proliferation of malicious or inappropriate volumetric content, underscoring the need for proactive measures to mitigate such risks.
Appendix A Additional Quantitative Evaluation Results
We report the detailed quantitative evaluation results, including storage size and rendering fidelity metrics (PSNR, SSIM, LPIPS), for each scene of the N3V and Technicolor datasets in Table 4. The most notable highlight is the substantial model size reduction achieved by our proposed method across all scenes, ranging from 79.3% at the highest fidelity setting (Level 6) to 99.3% at the most aggressive compression level (Level 1). Although aggressive compression (lower levels) occasionally leads to significant drops in rendering fidelity, our framework provides multiple trade-off options to flexibly adapt according to deployment scenarios and resource constraints. These results indicate promising directions for future work, focusing on enhancing reconstruction fidelity at lower compression levels to further broaden applicability.
Table 4: Quantitative results averaged over the N3V and Technicolor datasets. We compare the proposed method (L1-L6) against the Ex4DGS ex4dgs using standard quality metrics and model size.
[N3V] Coffee Martini [N3V] Cook Spinach
SSIM PSNR LPIPS Size(MB) SSIM PSNR LPIPS Size(MB)
Ex4DGS ex4dgs 0.918 28.79 0.070 129.77 Ex4DGS ex4dgs 0.956 33.24 0.042 118.38
Ours [$\mathbb{L}$6] 0.897 26.17 0.091 15.29 (-88.2%) Ours [$\mathbb{L}$6] 0.942 30.69 0.064 10.44 (-91.2%)
Ours [$\mathbb{L}$5] 0.894 25.92 0.098 10.60 (-91.8%) Ours [$\mathbb{L}$5] 0.939 30.50 0.069 6.75 (-94.3%)
Ours [$\mathbb{L}$4] 0.892 25.82 0.101 7.40 (-94.3%) Ours [$\mathbb{L}$4] 0.935 30.11 0.075 4.50 (-96.2%)
Ours [$\mathbb{L}$3] 0.888 25.74 0.108 5.21 (-96.0%) Ours [$\mathbb{L}$3] 0.930 29.55 0.083 3.12 (-97.4%)
Ours [$\mathbb{L}$2] 0.882 25.48 0.119 3.48 (-97.3%) Ours [$\mathbb{L}$2] 0.922 28.99 0.099 2.06 (-98.3%)
Ours [$\mathbb{L}$1] 0.868 24.97 0.144 1.92 (-98.5%) Ours [$\mathbb{L}$1] 0.910 28.38 0.121 1.15 (-99.0%)
[N3V] Cut Roasted Beef [N3V] Flame Salmon
SSIM PSNR LPIPS Size(MB) SSIM PSNR LPIPS Size(MB)
Ex4DGS ex4dgs 0.958 33.73 0.040 122.63 Ex4DGS ex4dgs 0.926 29.29 0.066 128.11
Ours [$\mathbb{L}$6] 0.943 31.45 0.064 10.04 (-91.8%) Ours [$\mathbb{L}$6] 0.910 26.58 0.079 13.43 (-89.5%)
Ours [$\mathbb{L}$5] 0.940 30.86 0.068 6.59 (-94.6%) Ours [$\mathbb{L}$5] 0.908 26.35 0.082 8.97 (-93.0%)
Ours [$\mathbb{L}$4] 0.936 30.23 0.074 4.42 (-96.4%) Ours [$\mathbb{L}$4] 0.905 25.92 0.088 6.07 (-95.3%)
Ours [$\mathbb{L}$3] 0.931 29.82 0.082 3.03 (-97.5%) Ours [$\mathbb{L}$3] 0.902 25.68 0.095 4.18 (-96.7%)
Ours [$\mathbb{L}$2] 0.924 29.04 0.093 2.01 (-98.4%) Ours [$\mathbb{L}$2] 0.893 24.87 0.107 2.77 (-97.8%)
Ours [$\mathbb{L}$1] 0.911 28.02 0.122 1.12 (-99.1%) Ours [$\mathbb{L}$1] 0.875 24.16 0.138 1.52 (-98.8%)
[N3V] Flame Steak [N3V] Sear Steak
SSIM PSNR LPIPS Size(MB) SSIM PSNR LPIPS Size(MB)
Ex4DGS ex4dgs 0.963 33.91 0.034 101.18 Ex4DGS ex4dgs 0.960 33.69 0.035 93.69
Ours [$\mathbb{L}$6] 0.950 31.34 0.059 8.49 (-91.6%) Ours [$\mathbb{L}$6] 0.950 31.70 0.056 8.66 (-90.8%)
Ours [$\mathbb{L}$5] 0.947 30.86 0.063 5.41 (-94.7%) Ours [$\mathbb{L}$5] 0.948 31.22 0.061 5.43 (-94.2%)
Ours [$\mathbb{L}$4] 0.943 30.38 0.070 3.56 (-96.5%) Ours [$\mathbb{L}$4] 0.944 30.62 0.068 3.58 (-96.2%)
Ours [$\mathbb{L}$3] 0.938 29.95 0.080 2.40 (-97.6%) Ours [$\mathbb{L}$3] 0.938 29.95 0.077 2.48 (-97.3%)
Ours [$\mathbb{L}$2] 0.931 29.13 0.093 1.58 (-98.4%) Ours [$\mathbb{L}$2] 0.931 29.34 0.093 1.64 (-98.3%)
Ours [$\mathbb{L}$1] 0.918 28.33 0.122 0.89 (-99.1%) Ours [$\mathbb{L}$1] 0.917 28.39 0.122 0.94 (-99.0%)
[Technicolor] Train
SSIM PSNR LPIPS Size(MB)
Ex4DGS ex4dgs 0.928 31.37 0.055 216.17
Ours [$\mathbb{L}$6] 0.931 31.55 0.040 33.19 (-84.6%)
Ours [$\mathbb{L}$5] 0.928 31.38 0.045 21.76 (-89.9%)
Ours [$\mathbb{L}$4] 0.913 30.39 0.060 12.73 (-94.1%)
Ours [$\mathbb{L}$3] 0.889 29.86 0.092 7.61 (-96.5%)
Ours [$\mathbb{L}$2] 0.845 28.44 0.146 4.55 (-97.9%)
Ours [$\mathbb{L}$1] 0.765 26.32 0.247 2.32 (-98.9%)
[Technicolor] Birthday [Technicolor] Fabien
SSIM PSNR LPIPS Size(MB) SSIM PSNR LPIPS Size(MB)
Ex4DGS ex4dgs 0.942 32.38 0.044 162.38 Ex4DGS ex4dgs 0.897 35.38 0.123 83.20
Ours [$\mathbb{L}$6] 0.932 31.33 0.046 33.62 (-79.3%) Ours [$\mathbb{L}$6] 0.871 32.90 0.168 5.89 (-92.9%)
Ours [$\mathbb{L}$5] 0.932 31.33 0.047 25.47 (-84.3%) Ours [$\mathbb{L}$5] 0.867 32.74 0.182 2.86 (-96.6%)
Ours [$\mathbb{L}$4] 0.929 31.08 0.051 18.50 (-88.6%) Ours [$\mathbb{L}$4] 0.861 32.60 0.200 1.77 (-97.9%)
Ours [$\mathbb{L}$3] 0.924 31.00 0.057 13.59 (-91.6%) Ours [$\mathbb{L}$3] 0.856 32.06 0.214 1.27 (-98.5%)
Ours [$\mathbb{L}$2] 0.916 30.56 0.068 9.34 (-94.2%) Ours [$\mathbb{L}$2] 0.850 31.68 0.228 0.90 (-98.9%)
Ours [$\mathbb{L}$1] 0.896 29.46 0.092 5.39 (-96.7%) Ours [$\mathbb{L}$1] 0.840 30.44 0.254 0.58 (-99.3%)
[Technicolor] Painter [Technicolor] Theater
SSIM PSNR LPIPS Size(MB) SSIM PSNR LPIPS Size(MB)
Ex4DGS ex4dgs 0.932 36.73 0.091 106.30 Ex4DGS ex4dgs 0.883 31.84 0.129 132.89
Ours [$\mathbb{L}$6] 0.914 35.05 0.110 9.19 (-91.4%) Ours [$\mathbb{L}$6] 0.871 30.16 0.138 16.27 (-87.8%)
Ours [$\mathbb{L}$5] 0.905 34.00 0.122 5.55 (-94.8%) Ours [$\mathbb{L}$5] 0.869 30.12 0.150 9.31 (-93.0%)
Ours [$\mathbb{L}$4] 0.891 33.29 0.142 3.47 (-96.7%) Ours [$\mathbb{L}$4] 0.858 29.70 0.172 5.58 (-95.8%)
Ours [$\mathbb{L}$3] 0.873 32.20 0.168 2.30 (-97.8%) Ours [$\mathbb{L}$3] 0.849 29.42 0.191 3.57 (-97.3%)
Ours [$\mathbb{L}$2] 0.848 30.96 0.210 1.47 (-98.6%) Ours [$\mathbb{L}$2] 0.833 29.00 0.221 2.19 (-98.4%)
Ours [$\mathbb{L}$1] 0.804 28.93 0.293 0.79 (-99.3%) Ours [$\mathbb{L}$1] 0.804 27.84 0.274 1.16 (-99.1%)
Appendix B Rate-Distortion Performance for Each Scene
In addition to the numerical results provided in Appendix A, we illustrate detailed per-scene rate-distortion (RD) curves for both datasets. Specifically, Figures 6 and 7 show RD curves for the N3V and Technicolor datasets, respectively. These figures visually reaffirm the notable size reductions achieved by our method, consistently demonstrating drastic reductions in model sizes. A closer examination reveals that fidelity degradation for the Technicolor dataset is relatively moderate compared to N3V, likely due to the simpler and less intricate motions inherent in Technicolor scenes. Conversely, N3V contains complex and detailed actions, such as precise tool movements or dynamic interactions (e.g., operating kitchen tongs or using a torch to sear meat), making compression-induced fidelity degradation more pronounced. This observation highlights opportunities for further enhancing compression methods, specifically targeting more effective exploitation of detailed volumetric motion information. Another noteworthy finding appears in the Technicolor dataset’s “Train” scene, where certain compression levels achieve superior rendering fidelity compared to the baseline Ex4DGS method. This result indicates the potential benefit of model simplicity, characterized by fewer points and lower entropy, which can unexpectedly enhance view-synthesis quality by effectively reducing model redundancy.
Figure 6: Per-scene RD performance comparison between our method (Levels 1-6) and the Ex4DGS ex4dgs on the Neural 3D Video (N3V) dataset.
Figure 7: Per-scene RD performance comparison between our method (Levels 1-6) and the Ex4DGS ex4dgs on the Technicolor dataset.
Appendix C Analysis on the Number of Pruned Points
To verify that the proposed method effectively reduces redundancy in both dynamic ($\mathcal{G}_{d}$) and static ($\mathcal{G}_{s}$) components, we examine the reduction ratios across different compression levels, as reported in Table 5. The results indicate that at the highest compression setting (Level 1), the reduction rates for dynamic and static points are similarly significant. However, at the lowest compression level (Level 6), the reduction ratio of dynamic points consistently surpasses that of static points. This suggests that when fidelity is prioritized, our method tends to prune dynamic points more aggressively, likely due to the smaller spatial coverage of dynamic points compared to static points. Nonetheless, given that dynamic points typically correspond to salient foreground elements, overly aggressive pruning of dynamic points may adversely impact overall visual quality. This highlights an avenue for future improvement: dynamically adjusting the pruning strategy—for example, by assigning separate weighting coefficients to static and dynamic points within the pruning loss—to better preserve visually salient dynamic components while maintaining high compression efficiency.
Table 5: Reduction in static ($\mathcal{G}_{s}$) and dynamic ($\mathcal{G}_{d}$) point counts for Technicolor scenes.
$\mathcal{G}_{s}$ $\mathcal{G}_{d}$ $\mathcal{G}_{s}$ $\mathcal{G}_{d}$
Birthday - 705377 1028398 Fabien - 87284 259209
1 62457 (-91.1%) 53578 (-94.8%) 1 7010 (-92.0%) 2455 (-99.1%)
6 419283 (-40.6%) 344054 (-66.5%) 6 47383 (-45.7%) 60256 (-76.8%)
Train - 1810464 241302 Painter - 103811 463955
1 57078 (-96.8%) 10013 (-95.9%) 1 4342 (-95.8%) 6307 (-98.6%)
6 1221140 (-32.6%) 57536 (-76.2%) 6 58010 (-44.1%) 103793(-77.6%)
Appendix D Visualization of Static and Dynamic Points
We visualize the static and dynamic components separately across different compression levels to further investigate the impact of our proposed compression method, as shown in Figures 8, 9 and 10. Consistent with the quantitative findings discussed in Appendix C, our method tends to prune dynamic points more aggressively than static points, despite dynamic regions generally holding higher visual saliency. Particularly noticeable in the highly compressed results (Level 1) shown in Figures 8 and 9, rapidly moving objects often exhibit noticeable local blur due to aggressive dynamic point pruning. Nevertheless, it’s worth emphasizing that even at moderate compression (Level 6), our method successfully achieves substantial storage savings (approximately 90% on average) while preserving rendering quality across most scenes without significant visual artifacts. This demonstrates our method’s practical balance between storage efficiency and visual fidelity.
Figure 8: Visualization of static and dynamic points on N3V dataset. From top to bottom: Coffee Martini, Cook Spinach, and Cut Roasted Beef.
Figure 9: Visualization of static and dynamic points on N3V dataset. From top to bottom: Flame Salmon, Flame Steak, and Sear Steak.
Figure 10: Visualization of static and dynamic points on Technicolor dataset. From top to bottom: Fabien, Birthday, Theater, Train, and Painter.
Appendix E Analysis of Compressed Representation Size
This appendix details the storage allocation (in MB and percentage) for individual components of our RD-optimized model across various scenes from the Technicolor Table 7 and N3V Table 6 datasets. Results are shown for Low Rate (Level 1) and High Rate (Level 6) settings. These tables provide a granular view of bit distribution for key components such as sorting indices for the masked SH parameters (sort idx), entropy-coded indexes (Indexes), masked wavelet coefficients $\mathbf{F}_{\text{masked}}$, dynamic position displacements ($\boldsymbol{\mu}_{disp}$), static position ($\boldsymbol{\mu}_{0}$), codebooks, logits, and opacity variance parameters ($\beta_{*}^{o}$).
Across both datasets, it can be observed that components related to the explicit representation of Gaussians and their dynamic properties (e.g., $\mathbf{F}_{\text{masked}}$, Indexes, sort idx) generally constitute a significant portion of the bitstream, especially at higher rates where more detail is preserved. The relative proportions vary across scenes, reflecting differences in scene complexity and motion characteristics.
Table 6: Storage allocation (MB and percentage) for components in our RD-optimized 4DGS model. N3V Scenes at Low ($\mathbb{L}$1) and High ($\mathbb{L}$6) rate settings.
Coffee Martini Cook Spinach
Low Rate [$\mathbb{L}$1] High Rate [$\mathbb{L}$6] Low Rate [$\mathbb{L}$1] High Rate [$\mathbb{L}$6]
Component Rate (MB) Proportion (%) Rate (MB) Proportion (%) Component Rate (MB) Proportion (%) Rate (MB) Proportion (%)
$\mathbf{F}_{\text{masked}}$ 0.448 22.21% 4.574 28.53% $\mathbf{F}_{\text{masked}}$ 0.285 23.71% 3.699 33.78%
Indexes 0.413 20.46% 4.051 25.26% Codebooks 0.241 20.01% 2.869 26.20%
sort idx 0.319 15.81% 2.574 16.05% Indexes 0.234 19.41% 1.409 12.87%
Codebooks 0.284 14.08% 1.676 10.45% sort idx 0.151 12.50% 0.862 7.87%
$\boldsymbol{\mu}_{0}$ 0.214 10.62% 1.676 10.45% $\boldsymbol{\mu}_{0}$ 0.097 8.07% 0.851 7.77%
$\boldsymbol{\mu}_{disp}$ 0.214 10.62% 0.999 6.23% $\boldsymbol{\mu}_{disp}$ 0.097 8.07% 0.851 7.77%
Logits 0.091 4.53% 0.339 2.11% Logits 0.078 6.46% 0.274 2.50%
$\beta_{*}^{o}$ 0.033 1.66% 0.146 0.91% $\beta_{*}^{o}$ 0.021 1.78% 0.135 1.24%
Header 0.000 0.00% 0.000 0.00% Header 0.000 0.00% 0.000 0.00%
Total 1.923 - 15.292 - Total 1.149 - 10.444 -
Cut Roasted Beef Flame Salmon
Low Rate [$\mathbb{L}$1] High Rate [$\mathbb{L}$6] Low Rate [$\mathbb{L}$1] High Rate [$\mathbb{L}$6]
Component Rate (MB) Proportion (%) Rate (MB) Proportion (%) Component Rate (MB) Proportion (%) Rate (MB) Proportion (%)
$\mathbf{F}_{\text{masked}}$ 0.262 22.34% 3.396 32.25% $\mathbf{F}_{\text{masked}}$ 0.303 19.03% 3.809 27.05%
Codebooks 0.238 20.29% 2.699 25.62% Indexes 0.303 18.98% 3.453 24.51%
Indexes 0.222 18.88% 1.432 13.60% Codebooks 0.271 17.01% 2.339 16.61%
sort idx 0.154 13.08% 0.885 8.41% sort idx 0.256 16.06% 1.542 10.95%
$\boldsymbol{\mu}_{0}$ 0.101 8.58% 0.885 8.41% $\boldsymbol{\mu}_{0}$ 0.175 10.99% 1.542 10.95%
$\boldsymbol{\mu}_{disp}$ 0.101 8.58% 0.849 8.06% $\boldsymbol{\mu}_{disp}$ 0.175 10.99% 0.973 6.91%
Logits 0.077 6.57% 0.252 2.39% Logits 0.088 5.51% 0.282 2.01%
$\beta_{*}^{o}$ 0.020 1.68% 0.133 1.27% $\beta_{*}^{o}$ 0.023 1.43% 0.143 1.02%
Header 0.000 0.00% 0.000 0.00% Header 0.000 0.00% 0.000 0.00%
Total 1.120 - 10.044 - Total 1.521 - 13.432 -
Flame Steak Sear Steak
Low Rate [$\mathbb{L}$1] High Rate [$\mathbb{L}$6] Low Rate [$\mathbb{L}$1] High Rate [$\mathbb{L}$6]
Component Rate (MB) Proportion (%) Rate (MB) Proportion (%) Component Rate (MB) Proportion (%) Rate (MB) Proportion (%)
Codebooks 0.222 23.68% 2.680 30.10% Codebooks 0.225 22.82% 2.738 30.15%
$\mathbf{F}_{\text{masked}}$ 0.183 19.54% 2.200 24.71% $\mathbf{F}_{\text{masked}}$ 0.189 19.21% 2.278 25.08%
Indexes 0.164 17.49% 1.263 14.18% Indexes 0.171 17.31% 1.282 14.12%
sort idx 0.121 12.90% 0.834 9.36% sort idx 0.134 13.56% 0.830 9.14%
$\boldsymbol{\mu}_{0}$ 0.090 9.11% 0.810 8.91% $\boldsymbol{\mu}_{0}$ 0.081 8.60% 0.798 8.97%
$\boldsymbol{\mu}_{disp}$ 0.081 8.60% 0.798 8.97% $\boldsymbol{\mu}_{disp}$ 0.090 9.11% 0.810 8.91%
Logits 0.072 7.70% 0.199 2.23% Logits 0.073 7.43% 0.203 2.24%
$\beta_{*}^{o}$ 0.014 1.47% 0.132 1.48% $\beta_{*}^{o}$ 0.014 1.45% 0.131 1.44%
Header 0.000 0.01% 0.000 0.00% Header 0.000 0.01% 0.000 0.00%
Total 0.895 - 8.491 - Total 0.941 - 8.661 -
Table 7: Storage allocation (MB and percentage) for components in our RD-optimized 4DGS model. Technicolor Scenes at Low ($\mathbb{L}$1) and High ($\mathbb{L}$6) rate settings.
Train Theater
Low Rate [$\mathbb{L}$1] High Rate [$\mathbb{L}$6] Low Rate [$\mathbb{L}$1] High Rate [$\mathbb{L}$6]
Component Rate (MB) Ratio (%) Rate (MB) Ratio (%) Component Rate (MB) Ratio (%) Rate (MB) Ratio (%)
sort idx 0.537 22.08% 10.230 29.40% $\mathbf{F}_{\text{masked}}$ 0.242 19.86% 5.660 33.17%
Indexes 0.458 18.81% 7.327 21.06% Codebooks 0.231 18.96% 4.272 25.04%
$\mathbf{F}_{\text{masked}}$ 0.361 14.84% 7.327 21.06% Indexes 0.222 18.24% 2.698 15.81%
$\boldsymbol{\mu}_{disp}$ 0.343 14.09% 6.394 18.37% sort idx 0.189 15.51% 1.258 7.37%
$\boldsymbol{\mu}_{0}$ 0.343 14.09% 2.072 5.95% $\boldsymbol{\mu}_{disp}$ 0.101 8.33% 1.080 6.33%
Codebooks 0.230 9.46% 0.866 2.49% $\boldsymbol{\mu}_{0}$ 0.101 8.33% 1.080 6.33%
Logits 0.080 3.31% 0.461 1.32% Logits 0.077 6.33% 0.883 5.18%
$\beta_{*}^{o}$ 0.080 3.31% 0.123 0.35% $\beta_{*}^{o}$ 0.054 4.43% 0.130 0.76%
Header 0.000 0.00% 0.000 0.00% Header 0.000 0.00% 0.000 0.00%
Total 2.319 - 33.187 - Total 1.162 - 16.271 -
Painter Fabien
Low Rate [$\mathbb{L}$1] High Rate [$\mathbb{L}$6] Low Rate [$\mathbb{L}$1] High Rate [$\mathbb{L}$6]
Component Rate (MB) Ratio (%) Rate (MB) Ratio (%) Component Rate (MB) Ratio (%) Rate (MB) Ratio (%)
$\mathbf{F}_{\text{masked}}$ 0.227 27.48% 3.737 38.79% Codebooks 0.188 31.09% 2.169 35.11%
Codebooks 0.196 23.66% 2.499 25.94% $\mathbf{F}_{\text{masked}}$ 0.089 14.67% 1.516 24.53%
Indexes 0.150 18.12% 1.295 13.44% Indexes 0.084 13.89% 0.861 13.94%
sort idx 0.085 10.33% 0.831 8.62% sort idx 0.076 12.57% 0.482 7.81%
Logits 0.065 7.91% 0.472 4.90% Logits 0.063 10.48% 0.480 7.77%
$\beta_{*}^{o}$ 0.051 6.13% 0.348 3.62% $\boldsymbol{\mu}_{disp}$ 0.042 7.00% 0.285 4.61%
$\boldsymbol{\mu}_{disp}$ 0.026 3.18% 0.348 3.62% $\boldsymbol{\mu}_{0}$ 0.042 7.00% 0.285 4.61%
$\boldsymbol{\mu}_{0}$ 0.026 3.18% 0.104 1.08% $\beta_{*}^{o}$ 0.020 3.29% 0.100 1.62%
Header 0.000 0.01% 0.000 0.00% Header 0.000 0.01% 0.000 0.00%
Total 0.789 - 9.188 - Total 0.576 - 5.892 -
Birthday
Low Rate [$\mathbb{L}$1] High Rate [$\mathbb{L}$6]
Component Rate (MB) Ratio (%) Rate (MB) Ratio (%)
$\mathbf{F}_{\text{masked}}$ 1.929 34.13% 12.386 35.14%
Indexes 1.280 22.65% 8.057 22.86%
sort idx 0.929 16.43% 6.107 17.32%
$\beta_{*}^{o}$ 0.429 7.59% 2.753 7.81%
$\boldsymbol{\mu}_{disp}$ 0.375 6.63% 2.516 7.14%
$\boldsymbol{\mu}_{0}$ 0.375 6.63% 2.516 7.14%
Codebooks 0.250 4.43% 0.794 2.25%
Logits 0.085 1.50% 0.122 0.35%
Header 0.000 0.00% 0.000 0.00%
Total 5.391 - 33.617 -